# arXiv:2603.00989v2[cs.SE] 03 Mar 2026

### v2

@@ V-G Energy Measurement and Hardware Evaluation Gaps
VI Limitations and Threats to Validity
@@ VI-B Internal Validity
To minimize these risks, we followed clearly defined inclusion and exclusion criteria throughout the screening process. Any uncertain cases were carefully examined and discussed to ensure consistent decisions. We also used structured data extraction forms to systematically collect and classify information such as evaluation metrics, model size categories, prompt engineering techniques, and sustainability aspects. Only information explicitly stated in the papers was recorded. If details were not clearly reported, we did not make assumptions. For instance, when a study used GitHub Copilot without specifying the underlying LLM, it was categorized as “not specified.” Although these steps cannot completely remove all bias, they help maintain a systematic, transparent, and consistent review process.
VI-C External Validity
@@
VII Conclusions
LLMs are increasingly being integrated into software engineering workflows to assist developers with tasks such as code generation, completion, translation, and bug fixing. While significant research has focused on improving the functional correctness and productivity gains of these models, comparatively little attention has been paid to the sustainability of the code they produce. We conduct a systematic literature review to examine whether LLMs can generate sustainable code. We follow a structured process that includes comprehensive database searches and snowballing method, and applying clear inclusion and exclusion criteria in multiple screening stages. For the selected studies, we extract data using predefined forms to ensure consistency and accuracy. Our findings show that research in this area is still limited. Most studies focus on large language models, while small models receive little attention. Energy evaluation is mainly performed at the software level, and only a few studies use external hardware tools for more accurate measurement. In addition, research is largely limited to a small number of programming languages and general-purpose applications. There are no dedicated benchmarks specifically designed to measure the sustainability of LLM-generated code, and sustainability-aware fine-tuning remains an open area for future work. To advance this field, future research should develop dedicated sustainability benchmarks, adopt more accurate energy measurement techniques, and explore sustainability-aware training and fine-tuning strategies. Addressing these gaps is essential to ensure that the growing use of LLMs in software development aligns with long-term environmental sustainability goals.

### v3

@@ VII-G Energy Measurement and Hardware Evaluation Gaps
VIII Limitations and Threats to Validity
@@ VIII-B Internal Validity
To minimize these risks, we followed clearly defined inclusion and exclusion criteria throughout the screening process. Any uncertain cases were carefully examined and discussed to ensure consistent decisions. We also used structured data extraction forms to systematically collect and classify information such as evaluation metrics, model size categories, prompt engineering techniques, and sustainability aspects. Only information explicitly stated in the papers was recorded. If details were not clearly reported, we did not make assumptions. For instance, when a study used GitHub Copilot without specifying the underlying LLM, it was categorized as “not specified.” Although these steps cannot completely remove all bias, they help maintain a transparent, and consistent review process.
VIII-C External Validity
@@
IX Conclusions
LLMs are increasingly being integrated into software engineering workflows to assist developers with tasks such as code generation, completion, translation, and bug fixing. While significant research has focused on improving the functional correctness and productivity gains of these models, comparatively little attention has been paid to the sustainability of the code they produce. We conduct a systematic literature review to examine whether LLMs can generate sustainable code. We follow a structured process that includes comprehensive database searches and a snowballing method, and applying clear inclusion and exclusion criteria in multiple screening stages. For the selected studies, we extract data using predefined forms to ensure consistency and accuracy. Our findings show that research in this area is still limited. Most studies focus on large language models, while small models receive little attention. Energy evaluation is mainly performed at the software level, and only a few studies use external hardware tools for more accurate measurement. In addition, research is largely limited to a small number of programming languages and general-purpose applications. There are no dedicated benchmarks specifically designed to measure the sustainability of LLM-generated code, and sustainability-aware fine-tuning remains an open area for future work. To advance this field, future research should develop dedicated sustainability benchmarks, adopt more accurate energy measurement techniques, and explore sustainability-aware training and fine-tuning strategies. Addressing these gaps is essential to ensure that the growing use of LLMs in software development aligns with long-term environmental sustainability goals.